Hi,
I am trying to find a solution to the following equation: log 4x = log 8 + 1. My instructor told me to move it to be: log 4 x - log 8 = 1. He said then to use the properties to write the equation as a single log. The next step was to re-write into exponential form and solve. I have tried all the methods I have used in class (one to one method, etc.) but we have not done a problem exactly like this one. I have looked and looked and cannot find a problem like this and it is driving me crazy. If you could show me an example of a problem like this, it would be appreciated.
Thanks in advance!

Answer by Personal Trainer DC(4)

log 4x - log 8 = 1
The property he is referring to is this:
log (4x/8) = 1
log (x/2) = 1
So, 10^1 = x/2
10 = x /2
x = 20